# Linear colorings of graphs

### v1

@@ 1 Introduction
Treedepth is a graph parameter that has been studied under different names, such as vertex ranking number [Sch89], ordered coloring [KMS95], and elimination height [Pot88]. It plays a central role in the graph sparsity theory of Nešetřil and Ossona de Mendez [NOdM12] due to its strong ties with the notion of bounded expansion. Like several other concepts from sparsity theory, a convenient aspect of this parameter is that it has several equivalent characterizations, so one can choose the most relevant one depending on the context.
The definition of treedepth that we mostly consider in this work is via centered colorings. A centered coloring of a graph $G$ is an assignment of integers (referred to as colors) to its vertices such that in every connected subgraph of $G$ there is a vertex with a unique color, called center. Observe that every centered coloring is proper (i.e., adjacent vertices receive different colors). The minimum number of colors required in a centered coloring of $G$ is its centered chromatic number, denoted by $\operatorname{\chi_{\mathrm{cen}}}(G)$. As it turns out, this number is equal to the treedepth of $G$ (see [NOdM12]).
Kun, O’Brien, Pilipczuk, and Sullivan introduced in [KOPS21] the following relaxation of centered colorings. A linear coloring of a graph $G$ is an assignment of integers to its vertices such that in every path of $G$ there is a vertex with a unique color. The linear chromatic number of $G$ is the minimal number of colors required in a linear coloring of $G$, and we denote it by $\operatorname{\chi_{\mathrm{lin}}}(G)$. As paths of $G$ are connected subgraphs, every centered coloring is a linear coloring, so we always have $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant\operatorname{\chi_{\mathrm{cen}}}(G)$. Observe that, just like centered colorings, every linear coloring is also proper. The motivation of the authors of [KOPS21] for introducing the linear chromatic number is algorithmic. They suggested an approach towards improving the running time (in practice) of an algorithmic pipeline for classes of bounded expansion [DRR+19, DKT13, NdM08] and were able to reduce the question whether their algorithmic strategy is relevant to the following graph theory problem: how big can $\operatorname{\chi_{\mathrm{cen}}}$ be compared to $\operatorname{\chi_{\mathrm{lin}}}$? In the same paper they provided the following partial answer: there is a constant $c\leqslant 190$ such that for every graph $G$, $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant\operatorname{\chi_{\mathrm{lin}}}(G)^{c}\cdot(\log\operatorname{\chi_{\mathrm{lin}}}(G))^{\operatorname{\mathcal{O}}(1)}$.
Using the recent results of [CNP21] the upper bound on $c$ could be improved to 19 (as observed in [KOPS21]) and [BDH+22] further improved it to 10.
@@
On the other hand, the authors of [KOPS21] provided a sequence of $P_{5}$-free chordal graphs for which the ratio $\operatorname{\chi_{\mathrm{cen}}}(G)/\operatorname{\chi_{\mathrm{lin}}}(G)$ can be arbitrarily close to $2$. They conjectured that this could be the optimal gap between the two invariants.
@@
[[KOPS21]] For every graph $G$, $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant 2\operatorname{\chi_{\mathrm{lin}}}(G)$.
@@
[[KOPS21, Theorem 4]] Let $T$ be a tree of maximum degree $\Delta\geqslant 3$. Then $\operatorname{\chi_{\mathrm{cen}}}(T)\leqslant(\log\Delta)\cdot\operatorname{\chi_{\mathrm{lin}}}(T)$.
In the case of interval graphs, [KOPS21] provided the following improvement over the general bound. {theorem}[[KOPS21]] For every interval graph $G$, $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant\operatorname{\chi_{\mathrm{lin}}}(G)^{2}$.
@@
We confirm Section 1 in several restricted graph classes such as caterpillars (for which we show that the two parameters differ by at most one) and several graph classes for which we can show that linear and centered colorings coincide: $(P_{3}+P_{1})$-free graphs, (claw, net)-free graphs (which contain in particular the proper interval graphs), complete multipartite graphs, and complements of rook’s graphs. [^2] See Section 4 for definitions. This is done by a close inspection of the structure of the considered graph classes. Note that the result for proper interval graphs cannot be generalized to the class of interval graphs, as shown by the caterpillars. Let us also remark that if there is a constant $c>0$ such that $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant c\cdot\operatorname{\chi_{\mathrm{lin}}}(G)$ for every chordal graph, then $c\geqslant 2$ due to the aforementioned construction from [KOPS21]. Our results in this direction are summarized in Table 1.
@@
The fact the centered and the linear chromatic numbers coincide for the class of complete multipartite graphs follows from the analogous result for cographs due to Kun et al. [KOPS21]. However, for complete multipartite graphs, as well as for complements of rook’s graphs, we determine the exact values of these parameters.
We then turn to another aspect of the linear chromatic number: obstructions. The class of graphs with linear chromatic number at most some constant $k$ is closed under subgraphs, so we can define its obstruction set (with respect to the subgraph relation) as the set of subgraph-minimal elements that do not belong to the class. Obstruction sets of graph classes closed under a certain graph containment relation are interesting, especially when they have bounded size, as they provide characterizations in terms of forbidden substructures in the spirit of Kuratowski’s theorem and, in case of the subgraph relation, immediately lead to polynomial-time recognition algorithms. In the case of the centered chromatic number, obstruction sets can be defined similarly. They have been shown to have finite size by Dvořák, Giannopoulou, and Thilikos [DGT12] (lower bounds can be found in the same paper). We observe that this property also holds for obstructions for bounded linear chromatic number (Corollary 5). In [DGT12] the set of obstructions for centered chromatic number at most $k$ is given for $k\in\{1,2,3\}$. By revisiting their proof, we provide obstruction sets for linear chromatic number at most $k$, for $k\in\{1,2,3\}$.
We conclude with algorithmic aspects. It is known that computing the centered chromatic number is NP-complete [BGT98]. We show that this also holds for computing the linear chromatic number (Section 6). Finally we provide an FPT algorithm for deciding if a graph has linear coloring number at most $k$ (Section 6).
@@ 2.1 Notations and definitions
In this paper logarithms are binary. Unless stated otherwise, we use standard graph theory terminology. The clique number of a graph $G$ is the maximum order of a clique in $G$ and is denoted by $\omega(G)$. The complement $\overline{G}$ of $G$ is the graph obtained by reversing the adjacency relation. For a subset $X\subseteq V(G)$, we use $G-X$ to refer to the graph obtained from $G$ after the deletion of the vertices in $X$. A graph is subcubic if its maximum degree is at most $3$.
@@
A graph coloring is a function mapping the vertices of a graph to integers, refered to as colors. A coloring is linear if every path has a unique color and centered if every connected subgraph has a unique color. The minimum number of colors (i.e., the cardinality of the image) of a linear (resp. centered) coloring of a graph $G$ is its linear chromatic number (resp. centered chromatic number), and we denote it by $\operatorname{\chi_{\mathrm{lin}}}(G)$ (resp. $\operatorname{\chi_{\mathrm{cen}}}(G)$). As noted in the introduction, $\operatorname{\chi_{\mathrm{cen}}}$ is the parameter most often called treedepth (see [NOdM12, Proposition 6.6]).
@@ 2.2 Centered versus linear
It is well known that the centered chromatic number does not increase when taking minors (see for instance [NOdM12, Lemma 6.2]). {proposition} If $H$ is a minor of $G$, then $\operatorname{\chi_{\mathrm{cen}}}(H)\leqslant\operatorname{\chi_{\mathrm{cen}}}(G)$.
@@
By Section 2.2, Section 2.2, and Section 1 respectively, we have $\operatorname{\chi_{\mathrm{lin}}}(H)\leqslant\operatorname{\chi_{\mathrm{cen}}}(H)\leqslant\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant f(\operatorname{\chi_{\mathrm{lin}}}(G))$, where $f$ is a polynomial of degree $11$ whose existence is implied by Section 1. ∎
However, in general it is not true that linear chromatic number is monotone under taking minors (in fact, not even under taking topological minors), as shown by the following construction due to Jana Masaříková and Wojciech Nadara (personal communication, 2024). We note that Section 1, if true, would imply that $\operatorname{\chi_{\mathrm{lin}}}(H)\leqslant 2\cdot\operatorname{\chi_{\mathrm{lin}}}(G)$ for every minor $H$ of $G$.
@@
Let $G$ be the graph from Fig. 1 (left) and $\psi$ be the coloring of $G$ specified by the numbers on its vertices in the aforementioned figure. Let $H$ be the graph from Fig. 1 (right) and observe that $H$ is a topological minor of $G$ that can be obtained by contracting the edge $kc$. We show that $G$ and $H$ satisfy the statement of the proposition. First, we verify that $\psi$ is a linear coloring by checking all the paths in $G$. Towards a contradiction suppose there is a path $P$ that does not have a center. Then, $P$ must see one of the colors 1 and 2. Suppose it sees 1 but not 2. Then, in order to see color 1 twice, $P$ contains vertices $g$, $f$, and $a$ in order. It follows that $P$ contains a center of color 0 or 3, a contradiction. By symmetry, it remains to analyze the case when $P$ sees both 1 and 2. In order to see colors 1 and 2 twice, $P$ starts in $g$ and ends in $h$. It cannot use $e$ (otherwise $P=gfedh$, admitting $g$ as a center) so it starts with $gfa$ and ends with $cdh$. In the rest of the graph (i.e., $G[\{i,b,j,k\}]$) no component contains twice the color 00, so any way to complete the path contains a center of color 0, a contradiction.
@@
[…]
Now, we show that $\operatorname{\chi_{\mathrm{lin}}}(H)>4$. Suppose, for contradiction, that there exists a linear coloring $\varphi$ of $H$ using four colors, say $\{0,1,2,3\}$. Observe that any cycle in $H$ has at least two centers in any linear coloring (since otherwise removing a unique center would result in a path without a center), in particular let $1,0,2$ be colors of $a,b$, and $c$ in $\varphi$, respectively. We split the analysis according to pairs of unique colors in the cycle $C$ with vertices $a,b,c,d,e,f$. Observe that two centers cannot be adjacent in $C$, as otherwise the rest of $C$ is a four-vertex path, which needs at least three colors in any linear coloring. Therefore, if 0 is a unique color in $C$, neither 1 nor 2 is unique in $C$. Moreover, observe that also 3 is not a unique color in $C$ as otherwise the vertices from $C$ colored by 1 and 2 would create a four-vertex path in $H$. This is a contradiction, since the cycle has two unique colors. It follows that 0 is not a unique color in $C$.
[…]
On paths the two invariants coincide (because every connected subgraph of a path is a path) and their value is logarithmic in their length (see [NOdM12, Section 6.2]). {lemma} For every integer $n$, $\operatorname{\chi_{\mathrm{cen}}}(P_{n})=\operatorname{\chi_{\mathrm{lin}}}(P_{n})=\lceil\log(n+1)\rceil$.
@@
It is easy to check that coloring each vertex of $B_{k}$ with its distance to the root is a centered coloring, and it has $k$ colors, so $\operatorname{\chi_{\mathrm{cen}}}(B_{k})\leqslant k$. Conversely, by contradiction suppose that $\operatorname{\chi_{\mathrm{cen}}}(B_{k})\geqslant k$ fails for some value of $k>1$ and let us consider the minimum such value. So $B_{k}$ admits a centered coloring $\varphi$ with less than $k$ colors. From the definition there is a vertex $v$ to which $\varphi$ assigns a unique color. Observe that $B_{k}-v$ contains a copy of $B_{k-1}$, that we call $B$. As $\varphi(v)$ is unique, $\varphi$ assigns at most $k-2$ colors to $B$. By minimality of $k$, $B$ has a subgraph with no unique color by $\varphi$, but this is also a subgraph of $B_{k}$ with this property, a contradiction. ∎
@@
Since the complete binary tree $B_{k}$ is subcubic, if $G$ contains $B_{k}$ as a minor, then it contains a subdivision $B$ of $B_{k}$ as a subgraph (see [Die25, Proposition 1.7.3]). As subdividing edges does never decrease $\operatorname{\chi_{\mathrm{cen}}}$ (Section 2.2) and by Section 2.2, we infer that $\operatorname{\chi_{\mathrm{cen}}}(B)\geqslant k$. By Section 2.2, $\operatorname{\chi_{\mathrm{lin}}}(B)\geqslant k/\log 3$. By Section 2.2, we get the desired bound. ∎
@@
[[KOPS21]] $\lim\limits_{k\rightarrow\infty}\operatorname{\chi_{\mathrm{cen}}}(B_{k})/\operatorname{\chi_{\mathrm{lin}}}(B_{k})\geqslant\log 3$.

### v2

@@ 1 Introduction
Treedepth is a graph parameter that has been studied under different names, such as vertex ranking number [SCH89], ordered coloring [KMS95], and elimination height [POT88]. It plays a central role in the graph sparsity theory of Nešetřil and Ossona de Mendez [NO12] due to its strong ties with the notion of bounded expansion. Like several other concepts from sparsity theory, a convenient aspect of this parameter is that it has several equivalent characterizations, so one can choose the most relevant one depending on the context.
The definition of treedepth that we mostly consider in this work is via centered colorings. A centered coloring of a graph $G$ is an assignment of integers (referred to as colors) to its vertices such that in every connected subgraph of $G$ there is a vertex with a unique color, called center. Observe that every centered coloring is proper (i.e., adjacent vertices receive different colors). The minimum number of colors required in a centered coloring of $G$ is its centered chromatic number, denoted by $\operatorname{\chi_{\mathrm{cen}}}(G)$. As it turns out, this number is equal to the treedepth of $G$ (see [NO12]).
Kun, O’Brien, Pilipczuk, and Sullivan studied in [KOP+21] the following relaxation of centered colorings due to Felix Reidl and Fernando Sánchez-Villaamil. A linear coloring of a graph $G$ is an assignment of integers to its vertices such that in every path of $G$ there is a vertex with a unique color. The linear chromatic number of $G$ is the minimal number of colors required in a linear coloring of $G$, and we denote it by $\operatorname{\chi_{\mathrm{lin}}}(G)$. As paths of $G$ are connected subgraphs, every centered coloring is a linear coloring, so we always have $\operatorname{\chi_{\mathrm{lin}}}(G)\leqslant\operatorname{\chi_{\mathrm{cen}}}(G)$. Observe that, just like centered colorings, every linear coloring is also proper. The motivation of the authors of [KOP+21] for introducing the linear chromatic number is algorithmic. They suggested an approach towards improving the running time (in practice) of an algorithmic pipeline for classes of bounded expansion [DRR+19, DKT13, Nd08] and were able to reduce the question whether their algorithmic strategy is relevant to the following graph theory problem: how big can $\operatorname{\chi_{\mathrm{cen}}}$ be compared to $\operatorname{\chi_{\mathrm{lin}}}$? In the same paper they provided the following partial answer: there is a constant $c\leqslant 190$ such that for every graph $G$, $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant\operatorname{\chi_{\mathrm{lin}}}(G)^{c}\cdot(\log\operatorname{\chi_{\mathrm{lin}}}(G))^{\operatorname{\mathcal{O}}(1)}$.
Using the recent results of [CNP21] the upper bound on $c$ could be improved to 19 (as observed in [KOP+21]), and [BDH+22] further improved it to 10.
@@
On the other hand, the authors of [KOP+21] provided a sequence of $P_{5}$-free chordal graphs for which the ratio $\operatorname{\chi_{\mathrm{cen}}}(G)/\operatorname{\chi_{\mathrm{lin}}}(G)$ can be arbitrarily close to $2$. They conjectured that this could be the optimal gap between the two invariants.
@@
[[KOP+21]] For every graph $G$, $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant 2\operatorname{\chi_{\mathrm{lin}}}(G)$.
@@
[[KOP+21, Theorem 4]] Let $T$ be a tree of maximum degree $\Delta\geqslant 3$. Then $\operatorname{\chi_{\mathrm{cen}}}(T)\leqslant(\log\Delta)\cdot\operatorname{\chi_{\mathrm{lin}}}(T)$.
In the case of interval graphs, [KOP+21] provided the following improvement over the general bound. {theorem}[[KOP+21]] For every interval graph $G$, $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant\operatorname{\chi_{\mathrm{lin}}}(G)^{2}$.
@@
We confirm Section 1 in several restricted graph classes such as caterpillars (for which we show that the two parameters differ by at most one) and several graph classes for which we can show that linear and centered colorings coincide: $(P_{3}+P_{1})$-free graphs, (claw, net)-free graphs (which contain in particular the proper interval graphs), complete multipartite graphs, and complements of rook’s graphs. [^2] See Section 4 for definitions. This is done by a close inspection of the structure of the considered graph classes. Note that the result for proper interval graphs cannot be generalized to the class of interval graphs, as shown by the caterpillars. Let us also remark that if there is a constant $c>0$ such that $\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant c\cdot\operatorname{\chi_{\mathrm{lin}}}(G)$ for every chordal graph, then $c\geqslant 2$ due to the aforementioned construction from [KOP+21]. Our results in this direction are summarized in Table 1.
@@
The fact that the centered and the linear chromatic numbers coincide for the class of complete multipartite graphs follows from the analogous result for cographs due to Kun et al. [KOP+21]. However, for complete multipartite graphs, as well as for complements of rook’s graphs, we determine the exact values of these parameters.
We then turn to another aspect of the linear chromatic number: obstructions. The class of graphs with linear chromatic number at most some constant $k$ is closed under subgraphs, so we can define its obstruction set (with respect to the subgraph relation) as the set of subgraph-minimal elements that do not belong to the class. Obstruction sets of graph classes closed under a certain graph containment relation are interesting, especially when they have bounded size, as they provide characterizations in terms of forbidden substructures in the spirit of Kuratowski’s theorem and, in case of the subgraph relation, immediately lead to polynomial-time recognition algorithms. In the case of the centered chromatic number, obstruction sets can be defined similarly. They have been shown to have finite size by Dvořák, Giannopoulou, and Thilikos [DGT12] (lower bounds can be found in the same paper). We observe that this property also holds for obstructions for bounded linear chromatic number (Corollary 5). In [DGT12] the set of obstructions for centered chromatic number at most $k$ is given for $k\in\{1,2,3\}$. (See also [KÜH25] for complements about obstructions with respect to the induced subgraph relation.) By revisiting their proof, we provide obstruction sets for linear chromatic number at most $k$, for $k\in\{1,2,3\}$.
We conclude with algorithmic aspects. It is known that computing the centered chromatic number is NP-complete [BGT98a]. We show that this also holds for computing the linear chromatic number (Section 6). Finally, we provide an FPT algorithm for deciding if a graph has linear chromatic number at most $k$ (Section 6).
@@ 2.1 Notations and definitions
In this paper logarithms are binary. Unless stated otherwise, we use standard graph theory terminology. The clique number of a graph $G$ is the maximum order of a clique in $G$ and is denoted by $\omega(G)$. The complement $\overline{G}$ of $G$ is the graph obtained by reversing the adjacency relation. For a subset $X\subseteq V(G)$, we use GXG-X to refer to the graph obtained from $G$ after the deletion of the vertices in $X$. A graph is subcubic if its maximum degree is at most $3$.
@@
A graph coloring is a function mapping the vertices of a graph to integers, referred to as colors. A coloring is linear if every path has a unique color and centered if every connected subgraph has a unique color. The minimum number of colors (i.e., the cardinality of the image) of a linear (resp. centered) coloring of a graph $G$ is its linear chromatic number (resp. centered chromatic number), and we denote it by $\operatorname{\chi_{\mathrm{lin}}}(G)$ (resp. $\operatorname{\chi_{\mathrm{cen}}}(G)$). As noted in the introduction, $\operatorname{\chi_{\mathrm{cen}}}$ is the parameter most often called treedepth (see [NO12, Proposition 6.6]).
@@ 2.2 Centered versus linear
It is well known that the centered chromatic number does not increase when taking minors (see for instance [NO12, Lemma 6.2]). {proposition} If $H$ is a minor of $G$, then $\operatorname{\chi_{\mathrm{cen}}}(H)\leqslant\operatorname{\chi_{\mathrm{cen}}}(G)$.
@@
By Section 2.2, Section 2.2, and Section 1, respectively, we have $\operatorname{\chi_{\mathrm{lin}}}(H)\leqslant\operatorname{\chi_{\mathrm{cen}}}(H)\leqslant\operatorname{\chi_{\mathrm{cen}}}(G)\leqslant f(\operatorname{\chi_{\mathrm{lin}}}(G))$, where $f$ is a polynomial of degree $11$ whose existence is implied by Section 1. ∎
However, in general it is not true that the linear chromatic number is monotone under taking minors (in fact, not even under taking topological minors), as shown by the following construction due to Jana Masaříková and Wojciech Nadara (personal communication, 2024). We note that Section 1, if true, would imply that $\operatorname{\chi_{\mathrm{lin}}}(H)\leqslant 2\cdot\operatorname{\chi_{\mathrm{lin}}}(G)$ for every minor $H$ of $G$.
@@
Let $G$ be the graph from Fig. 1 (left) and $\psi$ be the coloring of $G$ specified by the numbers on its vertices in the aforementioned figure. Let $H$ be the graph from Fig. 1 (right) and observe that $H$ is a topological minor of $G$ that can be obtained by contracting the edge $kc$. We show that $G$ and $H$ satisfy the statement of the proposition. First, we verify that $\psi$ is a linear coloring by checking all the paths in $G$. Towards a contradiction suppose there is a path $P$ that does not have a center. Then, $P$ must see one of the colors 1 and 2. Suppose it sees 1 but not 2. Then, in order to see color 1 twice, $P$ contains vertices $g$, $f$, and $a$ in order. It follows that $P$ contains a center of color 0 or 3, a contradiction. By symmetry, it remains to analyze the case when $P$ sees both 1 and 2. In order to see colors 1 and 2 twice, $P$ starts in $g$ and ends in $h$. It cannot use $e$ (otherwise $P=gfedh$, admitting $g$ as a center) so it starts with gfagfa and ends with cdhcdh. In the rest of the graph (i.e., $G[\{i,b,j,k\}]$) no component contains twice the color 0, so any way to complete the path contains a center of color 0, a contradiction.
@@
Note that three colors are not enough for a linear coloring as $G$ contains a path on 9 vertices (for instance, $\mathit{gfedcjabi}$), which requires $4$ colors, as we will note in Section 2.2. We conclude that $\operatorname{\chi_{\mathrm{lin}}}(G)=4$.
Now, we show that $\operatorname{\chi_{\mathrm{lin}}}(H)>4$. Suppose, for contradiction, that there exists a linear coloring $\varphi$ of $H$ using four colors, say $\{0,1,2,3\}$. Observe that any cycle in $H$ has at least two centers in any linear coloring (since otherwise removing a unique center would result in a path without a center), in particular let $1,0,2$ be colors of $a,b$, and $c$ in $\varphi$, respectively. We split the analysis according to pairs of unique colors in the cycle $C$ with vertices $a,b,c,d,e,f$. Observe that two centers cannot be adjacent in $C$, as otherwise the rest of $C$ is a four-vertex path, which needs at least three colors in any linear coloring. Therefore, if 0 is a unique color in $C$, neither 1 nor 2 is unique in $C$. Moreover, observe that 3 also is not a unique color in $C$ as otherwise the vertices from $C$ colored by 1 and 2 would create a four-vertex path in $H$. This is a contradiction, since the cycle has two unique colors. It follows that 0 is not a unique color in $C$.
Assume for now that colors 1 and 2 are unique in $C$. It follows that $\varphi(f)=\varphi(d)=3$ and $\varphi(e)=0$ or vice versa. In both cases, $\varphi(g),\varphi(h)\in\{1,2\}$ as any of them creates a four-vertex path with vertices $f,e,d$. Consider paths $\mathit{bcdefg}$ and $\mathit{bafedh}$ in the case where $\varphi(e)=0$ and paths $\mathit{bcdh}$ and $\mathit{bafg}$ in the case where $\varphi(e)=3$. Together with the previous argument, these paths force $g$ and $h$ to have different colors in $\{1,2\}$. However, the path $\mathit{gfacdh}$ has no center. It follows that at least one of the colors 1 and 2 is not unique in $C$.
Finally, assume that 1 and 3 are unique colors in $C$. As 2, and also 0, are not unique in $C$ by the previous analysis, we have that the vertices $f,e$, and $d$ have different colors in $\{0,2,3\}$. We now determine colors of $i$ and $j$. As $\varphi(i)\neq\varphi(b)$ and path $\mathit{fedcbi}$ enforces $\varphi(i)\neq 3$, we have $\varphi(i)\in\{1,2\}$. Hence, $\varphi(j)=3$ by considering path $\mathit{ibaj}$ or $\mathit{ibcj}$ and triangle $\mathit{acj}$. Moreover, $\varphi(i)=2$, by path $\mathit{jcdefabi}$. As $\varphi(d)\neq\varphi(c)$ and $\varphi(d)\neq 0$ by path $\mathit{ibcd}$, $\varphi(d)=3$. It follows $\varphi(e)=0$ by path $\mathit{jcde}$, and then $\varphi(f)=2$. Consider the color of vertex $g$: $\varphi(g)\neq 0$ by path $\mathit{jcdefg}$, $\varphi(g)\neq 1$ by path $\mathit{dcjafg}$, $\varphi(g)\neq\varphi(f)=2$, and $\varphi(g)\neq 3$ by path $\mathit{bcdefg}$. This contradicts that $\varphi$ is a linear coloring of $H$ with unique colors 1 and 3 in $C$, but also that $\varphi$ is a linear coloring of $H$ as we consider all cases up to the symmetry of $1$ and $2$. This shows that $\operatorname{\chi_{\mathrm{lin}}}(H)>4$. ∎
On paths, the two invariants coincide (because every connected subgraph of a path is a path) and their value is logarithmic in their length (see [NO12, Section 6.2]). {lemma} For every integer $n$, $\operatorname{\chi_{\mathrm{cen}}}(P_{n})=\operatorname{\chi_{\mathrm{lin}}}(P_{n})=\lceil\log(n+1)\rceil$.
@@
It is easy to check that coloring each vertex of $B_{k}$ with its distance to the root is a centered coloring, and it has $k$ colors, so $\operatorname{\chi_{\mathrm{cen}}}(B_{k})\leqslant k$. Conversely, by contradiction suppose that $\operatorname{\chi_{\mathrm{cen}}}(B_{k})\geqslant k$ fails for some value of $k>1$ and let us consider the minimum such value. So $B_{k}$ admits a centered coloring $\varphi$ with less than $k$ colors. From the definition there is a vertex $v$ to which $\varphi$ assigns a unique color. Observe that $B_{k}-v$ contains a copy of $B_{k-1}$, that we call $B$. As $\varphi(v)$ is unique, $\varphi$ assigns at most k2k-2 colors to $B$. By minimality of $k$, $B$ has a subgraph with no unique color by $\varphi$, but this is also a subgraph of $B_{k}$ with this property, a contradiction. ∎
@@
Since the complete binary tree $B_{k}$ is subcubic, if $G$ contains $B_{k}$ as a minor, then it contains a subdivision $B$ of $B_{k}$ as a subgraph (see [DIE02, Proposition 1.7.3]). As subdividing edges never decreases $\operatorname{\chi_{\mathrm{cen}}}$ (Section 2.2), and by Section 2.2, we infer that $\operatorname{\chi_{\mathrm{cen}}}(B)\geqslant k$. By Section 2.2, $\operatorname{\chi_{\mathrm{lin}}}(B)\geqslant k/\log 3$. By Section 2.2, we get the desired bound. ∎
@@
[[KOP+21]] $\lim\limits_{k\rightarrow\infty}\operatorname{\chi_{\mathrm{cen}}}(B_{k})/\operatorname{\chi_{\mathrm{lin}}}(B_{k})\geqslant\log 3$.